An ideal gas is enclosed in a cylinder with a movable piston on top of it. The piston has a mass of 8,000 g and an area of 5.00 cm2 and is free to slide up and down, keeping the pressure of the gas constant.
(a) How much work is done on the gas as the temperature of 0.180 mol of the gas is raised from 15.0°C to 350°C? 
 J 

Grade:12

1 Answers

Rituraj Tiwari
3 years ago
First law of thermodynamics gives us the relation.
q=ΔU+w
w=q−ΔU
For a constant pressure process,
q=nCp​ΔT
ΔU=nCv​ΔT(always)
Hence,w=n(Cp​−Cv​)ΔT=nRΔT=0.2×8.314×(300−20)=466J